OpenStudy (anonymous):

logic: x - 1 >= 0 x > 0 i dont understand when i can remove the equal sign, can somebody explain it to me?

OpenStudy (anonymous):

and what about \[x+1\ge0\] for example

OpenStudy (turingtest):

those are two different statements what is the difference between\[x\ge0\]and\[x>0\]is basically your question?

OpenStudy (anonymous):

or what about \[x+1\le0\] or \[x-1\le0\]

OpenStudy (turingtest):

one question at a time please jessica, you are asking different kinds of questions

OpenStudy (anonymous):

i understand what the sign means.. but like my example.. we gonna take this on: x -1 >= 0 this is the same as: x > 0

OpenStudy (anonymous):

but how?

OpenStudy (turingtest):

no it is not take it one step at a time what do\[x>0,x\ge0,x<0\x\le0\]all mean

OpenStudy (turingtest):

\[x-1\ge0\implies x\ge1\]which is not the same as \[x>0\]what gives you this impression?

OpenStudy (turingtest):

i mean, what makes you think they are the same?

OpenStudy (anonymous):

look better the equal sign is now gone..

OpenStudy (turingtest):

ok so let me get this straight: your question is "how is \(x-1\ge0\) the same as \(x>0\) ?" that answer is: it's not.

OpenStudy (anonymous):

never mind then, this is what my teacher said

OpenStudy (turingtest):

look, they are not the same, but one \(implies\) the other

OpenStudy (turingtest):

if \(x-1\ge0\) then \(x\ge1\) from thinking about the number line, it is obvious that if a number is \(\text{at least 1}\) then is must always be greater than zero, so we can write\[x\ge1\implies x>0\]the arrow is read as the word "implies" the statement above is " since x is at least 1, that implies that x is more than zero" which makes sense, right?

OpenStudy (anonymous):

it does make sense, but im still confused a little bit

OpenStudy (turingtest):

we cannot write\[x\ge1\implies x\ge0\] because that is like saying "since x is at least 1, that implies that x is at least 0" but \(at\) \(least\) implies x it cannot be less than 1, so x=0 is not possible

OpenStudy (turingtest):

... a contradiction^ the fact that\[x\ge1\implies x>0\]is trivial "if the number is at least 1 it's more than zero" is a trivial statement we could just as easily have written\[x\ge1\implies x>\frac12\implies x>\frac14\implies x>-420,\text{etc.} \]because \[x\ge0\] implies that x is bigger than all those numbers I wrote above

OpenStudy (turingtest):

you could read the above line that is second tp last as "if x is at least 1, then it's more than 1/2, and 1/4, and -420" of course we could write any number less than one in those places

OpenStudy (anonymous):

oehh... i still find it difficult :S

OpenStudy (turingtest):

new ideas are always a little strange, but I promise you that if you meditate on it for a while it will seem kinda obvious I often think of \(\ge\) which is "greater-than or equal-to" as the phrase "at least" and \(\le\) which is "less-than or equal-to" as "at most so\[0\le x\le1\]means "x is at least 0 and at most 1" drawing number lines also helps in these situations too

OpenStudy (turingtest):

you tell me, what is the difference between\[x\ge0\]and\[x>0\]

OpenStudy (anonymous):

x>=0 is: 0,1,2,3,4.... x>0 is: 1,2,3,4....

OpenStudy (turingtest):

why are you using only integers? x>0 is 1/2, 3/10 as well but yeah, the equals sign means that number is \(included\) in the interval of x we are talking about without the equals sign the number itself is not included

OpenStudy (turingtest):

it is like the difference between the phrases "at least" and "more than"

OpenStudy (anonymous):

its just an habbit, i always use discreet :P

OpenStudy (turingtest):

\[x\ge0\]x is at least zero\[x>-5\]x is more than -5 -notice those statements do not contradict also notice that -5 could have been any number less than 0